OpenStudy (kaythxbai):

how many solutions does −6(x+7)=−4x−2 have?

OpenStudy (princeharryyy):

one solution

OpenStudy (princeharryyy):

-6x-42 = -4x-2 -2x = 42-2 -2x = 40 x = -20

OpenStudy (kaythxbai):

thanks! could you help me with this one too? 4960z+50−97z=−37z+49

OpenStudy (princeharryyy):

4960z-97z+37z = -50+49 4900z = -1 z= -1/4900

OpenStudy (kaythxbai):

so how many solutions does it have?

OpenStudy (princeharryyy):

one

OpenStudy (kaythxbai):

it says it was wrong

OpenStudy (princeharryyy):

what did u fill?

OpenStudy (kaythxbai):

i put one solution and it said it was wrong

OpenStudy (princeharryyy):

can u take snapshot the question?

OpenStudy (kaythxbai):

i cant im on laptop

OpenStudy (princeharryyy):

laptop allows screenshot

OpenStudy (kaythxbai):

60z+50−97z=−37z+49 thats how it exactly looks on the page, i copy-pasted it

OpenStudy (princeharryyy):

4960z+50−97z=−37z+49, this was ur equation earlier. LOL

OpenStudy (kaythxbai):

oh sorry! i might've copied the example XD

OpenStudy (princeharryyy):

try zero as your answer

OpenStudy (kaythxbai):

The answers i can put are no soultion, one soultion or infinate many soultion

OpenStudy (kaythxbai):

oh i just clicked no soultion and it worked

OpenStudy (princeharryyy):

zero solutions means no solution

OpenStudy (kaythxbai):

thanks ^^ could you help me with this last one?

OpenStudy (princeharryyy):

ok

OpenStudy (kaythxbai):

−5(z+1)=−2z+10

OpenStudy (princeharryyy):

one solution